OpenStudy (anonymous):

solve. x-7 = 3 ---- --- x+8 7

OpenStudy (anonymous):

can u re write this plaese

OpenStudy (anonymous):

x=10 x=15

OpenStudy (anonymous):

is this a fraction??

OpenStudy (anonymous):

ya ^^^^^^^^

myininaya (myininaya):

7(x-7)=3(x+8) 7x-49=3x+24 7x-3x=24+49 4x=73 x=73/4

OpenStudy (anonymous):

tina ann, do you know how to cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh well, myininaya already gave you the answer nvm

myininaya (myininaya):

sorry i stole your thunder dhatraditya

OpenStudy (anonymous):

lol